OpenStudy (anonymous):

Simplify the expression. x sqrt 8x^3 + sqrt2 + 4 sqrt2x If the simplified expression is written in standard form, what is the leading coefficient? PLEASE ANSWER WITH A NUMERICAL ANSWER

OpenStudy (anonymous):

Is this the question?\[x \sqrt {8x^3} + \sqrt2 + 4 \sqrt{2x}\]

OpenStudy (anonymous):

oops, no. It's x sqrt8x ^3 + xsqrt2+4sqrt2x

OpenStudy (anonymous):

I still can't tell where to start. Can you use the equation editor to help make it clear?

OpenStudy (anonymous):

Is this it? \[x \sqrt {8x^3} + \sqrt2x + 4 \sqrt{2x}\]

OpenStudy (anonymous):

\[x \sqrt{8x^3} + x \sqrt{2} + 4\sqrt{2x}\]

OpenStudy (anonymous):

Good. Here we go.\[x \sqrt {8x^3} + \sqrt2 x+ 4 \sqrt{2x}=2x^2\sqrt{2x}+\sqrt2x+4\sqrt{2x}\]

OpenStudy (anonymous):

The leading term in standard form is the first one (power of x is 5/2). The coefficient of that term is 2sqrt2, approximately 2.82

OpenStudy (anonymous):

Maybe one more step will make this clearer....\[2x^2\sqrt{2x}+ x \sqrt2 + 4 \sqrt{2x}=2\sqrt2x^2 \sqrt x+ x \sqrt2 + 4 \sqrt x\]